OpenStudy (anonymous):

3x+27=13x-83 = ? please help me lol.

OpenStudy (amoodarya):

27+83=13x-3x 110=10x so x=110/10 x=11

OpenStudy (amoodarya):

3x+27=13x-83=60

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

so much. lol